OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 5x - 3y - z if x = -2, y = 2, and z = -3 Medal?

OpenStudy (anonymous):

Would the answer be -19?

OpenStudy (anonymous):

no,,the answerr is -13

OpenStudy (anonymous):

see how... 5(-2)-3(2)-(-3)=-10-6+3=-16+3=-13

OpenStudy (anonymous):

How is that? 5*-2 would be -10. -3*2 would be -6. and you add -10 too -6, too -3 to get -19

OpenStudy (anonymous):

see it carefully...above..

OpenStudy (anonymous):

Oh. I see it now. Thanks